OpenStudy (anonymous):

Solve M = 2x + 3y for y.

OpenStudy (anonymous):

y=(M-2x)/3

OpenStudy (anonymous):

can you show steps

OpenStudy (anonymous):

you have to rearrange the formula so you subtract 2x from M so it's on the other side: M-2x=3y, now you have to divide by 3 to get y by itself and therefore get y= (M-2x)/3=y